Q1 ) The Ministry of public health considering a project  a the following cash flows: benefits of $214,726 in the first years and increase by 3,993 for through year 20,  disbenefits  of $49,852 in the first year and decreases by $1,148 through year 20, The project first cost $687,400 and M&O costs of $25,000 per year. If the discount rate is 10% and study period is 20 years the modified B/C ratio of the project  is closest to

Q2) A local government considering two projects with the following details: The project W has a benefits of $24,013 per year, disbenefits = $6,391 per year, and costs =$9,505per year. Project Z has a $38,515 with no disbenefits and with annual costs of $27,204 the incremental B/C ratio is closest to

Q3) Word health organization (WHO) plans to improve organic agriculture in poor countries of in the Midwestern Africa. One offers has the following details: a first cost of 1,128,251 with  M&O payments $42,611,  occurring  every fifth year, starting from year 5 through 20 years. The annual estimated benefits  is   $452,518 in year 1 and expected to increase by 8,812 starting from year 2 through the project period. However, local framers claims the the project is going to reduce revenue from  traditional  agriculture as follows: an amount  $84,604 in year 1 but decrease by 8,812 in year 2 through year 20. If the discount rate is 8%, the B/C ratio is  
Answer #1

i = 10%, t = 20 yrs

AW of benefits = 214726 + 3993*(A/G,10%,20) = 214726 + 3993 * 6.508075 = 240712.74

AW of disbenefits = 49852 - 1148*(A/G,10%,20) = 49852 - 1148*6.508075 = 42380.73

AW of cost = 687400*(A/P,10%,20) = 687400*0.117460 = 80742

Modified B/C ratio of project = (240712.74 - 42380.73 - 25000) / 80742

= 2.14674 = 2.15

Pls upload rest of the questions separately as per Chegg policy